OpenStudy (anonymous):

From the sum of (4x-3xy+2z-5) and (2z+3y-1+4x), subtract (5-4x+3y+2z) I need solution.

OpenStudy (anonymous):

Combine like terms. 4x + 4x = a 3xy = 3xy 3y = 3y 2z + 2z = b -1 - 5 = c a + 3xy + 3y + b - c